OpenStudy (anonymous):

Daily use of cigarettes by U.S. 12th graders is declining. In 1995, 21.6% of 12th graders smoked daily. This number decreased to 11.4% in 2008. Find the average rate of change in the percent of 12th graders who smoke from 1995 to 2008.

OpenStudy (anonymous):

Average rate of change is \[ \frac{f(b)-f(a)}{b-a} = \frac{11.4\%-21.6\%}{2008-1995} \]

OpenStudy (anonymous):

@wio I got -.784 but would I round?